OpenStudy (anonymous):

solve the following equation for x: -3(6x+5)=21

OpenStudy (solomonzelman):

\[-3(6x+5)=21\]\[-18x-15=21\][-18x-15=21\]\[~~~~~~~~~+15~~~+15\]\[-18x=36\]\[x=?\]

OpenStudy (anonymous):

x=-3*6x=-18x+-3*5=-15 so now you have... -18x+-15=21 x=21-(-15)or 21+15=36 x=36/-18 x=2 because -18*2=36